Firm Target (TGT) Walmart (WMT) Amazon (AMZN) ROA 7.68 4.892 2.83 EM 3.33 2.63 4.74 Based on the information in the table above, which firm has the highest debt to equity ratio? O Target O Walmart O Amazon Moving to another question will save this response.
Answer #1

Answer is 3 - Amazon

Walmart debt ratio = 1-(1/EM) = 1- (1/2.63) =61.97%

Target debt ratio = 1-(1/EM) = 1-(1/3.33) =69.99%

Amazon debt ratio = 1-(1/EM) = 1-(1/4.74) =78.90%